given the linear equation y = -7/4 x-1, find the y-coordinates of the points (-4, ), (0 ), and (4, ) please show all of your work. plot these points and graph the linear equation.
Answer by richwmiller(17219)

Those are very easy to plug in.
y=-7/4*(-4)-1
and
y=0-1
and
y=-7/4*4-1